OpenStudy (anonymous):

What is the value of sin 67° to the nearest ten-thousandth? What is the value of cos 34° to the nearest ten-thousandth? 4√6/√30 simplify by rationalizing the denominator

OpenStudy (anonymous):

do you want us to complete the first two without a calculator? or with?

OpenStudy (anonymous):

with

OpenStudy (anonymous):

sin(67) = 0.9205 (to the nearest ten-thousandth) cos(34) = 0.8290 (again rounded to the nearest ten-thousandth) \[\frac{4\sqrt{6}}{\sqrt{30}} = \frac{4\sqrt{3}\sqrt{2}}{\sqrt{3}\sqrt{10}} = \frac{4\sqrt{2}}{\sqrt{10}}\]\[\frac{4\sqrt{2}}{\sqrt{10}}=\frac{4\sqrt{2}}{\sqrt{2}\sqrt{5}} = \frac{4}{\sqrt{5}}\]\[\frac{4}{\sqrt{5}}= \frac{4}{\sqrt{5}}\times\frac{\sqrt{5}}{\sqrt{5}} = \frac{4\sqrt{5}}{5}\]

OpenStudy (anonymous):

you can split up radicals: \[\sqrt{ab} = \sqrt{a}\sqrt{b}\]

OpenStudy (anonymous):

thankyouu

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

could you help me with some of my other math PLEASE i dont understand it at all